A city when its operation is normal uses a supply of drinking water

containing 15 mg /L of chloride (Cl-), but during periods of high water demand,

mixes with groundwater containing 65 mg/L Cl-. The two waters mix

completely in a 1400 m3 tank and then sent to the distribution system. the

demand for many days has been ~500 m3/d, and all the water has come from

the main supply. So, the demand increases to 700 m3 /d, and the additional

200 m3/d is extracted from the groundwater supply. How long can

continue with this operating strategy before the water entering the

distribution system contains ≥ 20 mg/L Cl-? (Note: chloride is not reactive).


1
Expert's answer
2022-04-14T04:33:03-0400

 

15 mg/L = 15 g/m3 – normal supply concentration

65 mg/L = 65 g/m3 – ground water concentration

15 * 500 = 7500 g of Cl- - normal supply in one day

65 * 200 = 13000 g of Cl- - ground water in one day

Together: 7500 + 13000 = 20500 g

New concentration : 20500 / 700 = 29.3 g/m3 or 29.3 mg/L

So, in less than a day time the concentration is over the limit.